OpenStudy (anonymous):

e^x =3 PLEASE help me!!!

OpenStudy (anonymous):

take the natural log of both sides ln (e^x) = ln(3)

OpenStudy (anonymous):

this becomes x = ln(3)

OpenStudy (anonymous):

plug in calc and you got it